# SonarSplat: Novel View Synthesis of Imaging Sonar via Gaussian Splatting

### v1

@@ 3.3 Range/Azimuth Splatting Model
$\mu^{\prime}=K\mu_{s},\ \ \Sigma^{\prime}=J_{K}W\Sigma_{s}WJ_{K}^{T}$ (10)
where $J_{K}$ is the Jacobian of $K$, and $W$ is the sensor’s view matrix, similar to the camera splatting formulation [13].
To find the intensity of acoustic returns at a given range-azimuth bin ($r_{i},\theta_{j}$), we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range-azimuth image space that overlap with the range-azimuth bin ($r_{i},\theta_{j}$). The rasterization equation to obtain the sonar image for SonarSplat is similar to that of [13], with a few key differences. SonarSplat’s rasterization equation is given by:
$I_{U}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}\nu_{k}~{}o_{k}~{}T_{k}~{}G_{k}(r_{i},\theta_{j})$ (11)
where we introduce $\nu_{k}\in[0,1]$ as the acoustic reflectance of the Gaussian and $T_{k}$ as the transmittance from the sensor origin to the Gaussian along the range. The subscript indicates that $I_{U}$ is the unsaturated image, meaning that it does not account for the azimuth streaking that causes saturation across the image. We observe that the acoustic reflectance of a set of points depends on material properties and viewing angle, so $\nu_{k}$ uses spherical harmonics to encode view-dependent effects.
Figure 3: Azimuth streaking is characterized by black streaks (referred to by blue arrows) spanning the azimuth axis of a sonar image. (a) Sonar image of a metal frame structure that experiences azimuth streaking. (b) Sonar image of a concrete piling that has multiple azimuth streaks. Image brightness was enhanced for illustrative purposes. Images are polar (range/azimuth).
3.4 Modeling Azimuth Streaking
Azimuth streaking, as shown in Fig. 3, is a phenomena that is frequently observed in sonar images. It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to parallel at a specific range bin $r_{i}$ [29]. The azimuth streaking model found in [29] proposes to modify all the returns in a range interval if a certain percentage of returns exceeds a given threshold $\zeta$. The image with added azimuth streaks is given by:
$\hat{I}(r_{i},\cdot)=1-(1-I(r_{i},\cdot))^{2}$ (12)
where $\hat{I}(r_{i},\cdot)$ is the streaked image, and $I(r_{i},\cdot)$ is the vector of intensities at range interval $r_{i}$.
In our model, we assign each Gaussian a probability that it will cause an azimuth streak. During the splatting process, we accumulate this probability as a function of the Gaussian’s opacity, mean, and covariance. The benefit of defining this probability per-Gaussian is that it can be optimized to produce azimuth streaks that are multi-view consistent.
Let $p_{k}\in[0,1]$ be the probability that Gaussian $k$ will contribute to an azimuth streaking phenomena. We can then splat these per-Gaussian probabilities into a range-azimuth bin to obtain an azimuth streak probability image $P_{a}$:
$P_{a}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}p_{k}~{}o_{k}~{}T_{k}~{}G_{k}(r_{i},\theta_{j})$ (13)
@@ 3.4 Modeling Azimuth Streaking
3.4.1 Adaptive Gain Module
We introduce a novel adaptive gain term, $A$, to transform the azimuth streak probability image into receiver gains. We begin designing this gain term with a few observations. First, when no Gaussian in a range bin $r_{i}$ has a high probability of azimuth streaking, the gain should be unity. Secondly, suppose a single Gaussian $G_{k}$ exhibits a high probability of azimuth streaking across the range bin $r_{i}$. In that case, we wish to adaptively suppress the other Gaussians in $r_{i}$ and assign a high gain to $G_{k}$. Finally, if multiple Gaussians in $r_{i}$ have high probabilities of contributing to an azimuth streak, higher gain should be assigned to those Gaussians. To address these observations, the adaptive gain term is applied to each range-azimuth bin $(r_{i},\theta_{j})$:
$A(r_{i},\theta_{j})=P_{a}(r_{i},\theta_{j})M_{a}(r_{i})\frac{e^{\gamma\cdot P_{a}(r_{i},\theta_{j})}+1}{e^{\gamma}+1}+(1-M_{a}(r_{i}))$
where $\gamma$ is a scaling factor that dictates the steepness of the adaptive gain. This behavior aligns with the insight presented in [29]: if a certain percentage of returns exceeds a threshold, an operation is applied to all the values in a range interval. However, our proposed model differs because we do not restrict the gain to a fixed function (quadratic) but rather offer a family of curves for our splatting model to explore during optimization. Further discussion on the adaptive gain term can be found in the supplementary material.
The final image with the azimuth streaks is then computed by applying the gain $A(r_{i},\theta_{j})$ to its corresponding bin on $I_{U}(r_{i},\theta_{j})$:
@@ 3.4.1 Adaptive Gain Module
resulting in the final rendered sonar image $\hat{I}(r_{i},\theta_{j})$.
@@
Figure 4: We present selected images from our diverse set of 8 sequences used to evaluate of our method. We show RGB images of the environment for visualization, and captured range/azimuth sonar images. Our evaluation datasets focus on smaller objects and larger-scale structures such as the test-tank basin. Sonar images are polar (range/azimuth).
3.5 Model Optimization
We optimize for desired parameters using SGD and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$[13]. To properly optimize for the azimuth streaking probabilities $p_{k}$ of each Gaussian, we first identify where azimuth streaks occur by calculating the average intensity of the range interval $r_{i}$. Then, for the first $N_{s}$ iterations, we train only on pixels in range bins that exceed a certain average intensity threshold $\tau$. Importantly, $p_{k}$ is not optimized during this interval. Then, after $N_{s}$ iterations, we train only on pixels less than $\tau$. We do not optimize $\mu_{k},\Sigma_{k},o_{k}$, or $r_{k}$ during this interval. This way, we allow the Gaussians to initialize and fit the images before isolating and optimizing the azimuth streaking parameters.
We also encourage each Gaussian to have either high or low opacity via an opacity loss $\mathcal{L}_{o}$. This prevents a set of medium-opacity Gaussians from producing the same pixel intensity. Rather, the model is encouraged to use the reflectance parameter $r_{k}$ to represent low-intensity returns. We use the negative log-likelihood of two Laplacian distributions as in [15].
$\mathcal{L}_{o}(x)=|x||1-x|$ (17)
The final loss becomes:
$\mathcal{L}=\lambda_{l1}\mathcal{L}_{l1}+\lambda_{ssim}\mathcal{L}_{ssim}+\lambda_{o}\mathcal{L}_{o}$ (18)
@@ 3.5 Model Optimization
where $\lambda_{l1},\lambda_{ssim},\text{and}~{}\lambda_{o}$ are the weights for the L1, SSIM, and opacity losses respectively.
4 Results & Experiments
4.1 Datasets
We note that real, public datasets of posed sonar images suitable for novel view synthesis experiments do not exist. Rather, we collect sequences using our robotic platform based on a BlueRobotics BlueROV2 vehicle shown in Fig. 1. Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar capable of capturing data across various ranges with a 130 horizontal field of view and a 20 vertical field of view. We use this platform to collect 8 trajectories inspecting a range of objects and in both a test tank and river environment. Fig. 4 shows selected sequences for visualization. For detailed information about sonar parameters, sensor pose estimation, our platform, and dataset details, please refer to the supplemental material.
4.2 Baselines
For baselines, we compare to Neusis [30], Neusis-NGP [44], DSC [7], and ZSplat*. Note that ZSplat [32] was originally designed for camera-sonar fusion. Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with the sonar loss.
@@ 4.2 Baselines
Figure 5: We present qualitative novel view synthesis results from selected datasets and baselines. Note that SonarSplat produces superior sonar image synthesis compared to baselines and is better able to capture finer details in the environment. SonarSplat is also able to capture azimuth streaking phenomena, which no other baseline is able to do. We increase the brightness for all outputs by 25% for easier viewing. Note that Z-Splat* indicates Z-Splat trained only with the sonar loss. Images are cartesian.
De-Streaking Performance (ICV) ($\uparrow$)
Sequence Captured Image SonarSplat NVS De-Streaked Image
basin_empty 2.74 9.78 19.44
infra_360 2.89 34.30 44.18
piling_1 2.78 11.47 19.51
basin_infra 2.77 4.41 4.84
rock_360 2.86 25.00 44.47
piling_2 2.79 7.61 10.32
pole_river 3.11 2.69 2.78
rock_river 3.15 5.15 6.78
Average 2.89 12.55 19.04
Table 1: Quantitative results for azimuth streak removal. We report the inverse coefficient of variation (ICV) [37]. Higher is better ($\uparrow$), and best is in bold.
4.3 Novel View Synthesis
We train our method on an NVIDIA RTX 3090 GPU with 24 GB of vRAM. For detailed implementation details, please refer to the supplemental material. We report validation PSNR, SSIM, and LPIPS following [13, 21]. For our validation set, we take every 8th image from the dataset as introduced in [21]. We report SonarSplat’s performance compared to baselines in Tab. 2. We find that SonarSplat consistently synthesizes more realistic sonar images, quantified by its higher PSNR, SSIM, and lower LPIPS values. We also present qualitative results from selected sequences in Fig. 5. Note that these qualitative results are in cartesian space for interpretability. For extensive polar image results, please refer to the supplemental material. Finally, we explore the feasibility of SonarSplat for data synthesis applications by reporting the rendering speed in frames per second (FPS) of each method in Tab. 3. We calculate the FPS of each method on an NVIDIA A6000 GPU with 50 GB of vRAM. SonarSplat and ZSplat both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods. ZSplat is optimized for high rendering speed, which will be a focus of future work for SonarSplat.
@@ 4.3 Novel View Synthesis
Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Neusis [30] 14.43 0.03 0.66
Neusis-NGP [44] 19.43 0.33 0.55
DSC [7] 14.87 0.10 0.63
Z-Splat* [32] 8.57 0.03 0.68
SonarSplat (ours) 21.89 0.41 0.62
Table 2: Novel view synthesis performance on real world data. Note that ZSplat* is ZSplat with only the sonar loss activated for training. Best is shown in bold.
Figure 6: Qualitative results for azimuth streak removal capabilities of SonarSplat. Note that other baselines are unable to remove azimuth streaks from synthesized images since they do not learn the streaking parameters. Locations of azimuth streaks in the captured images are shown in blue arrows for illustrative purposes. SonarSplat is able to effectively restore sonar imagery that exhibits azimuth streaking. Images are cartesian.
Rendering Speed ($\uparrow$)
Method FPS
Neusis [30] 0.03
Neusis-NGP [44] 0.01
DSC [7] 0.03
Z-Splat* [32] 155.32
SonarSplat (ours) 9.46
Table 3: Rendering speed in frames per second (FPS) averaged across all validation images across all sequences. All results are reported on an NVIDIA RTX A6000 GPU. Note that ZSplat* is ZSplat with only sonar rendering. Best is shown in bold.
Figure 7: Qualitative results from 3D reconstruction of SonarSplat compared to baselines.
4.4 Azimuth Streak Removal
@@
Since SonarSplat learns probabilities of azimuth streaking for each Gaussian, we can undo the adaptive gain to render images without azimuth streaks present and recover suppressed returns. This vision task is most similar to image de-striping from the remote sensing literature. We present quantitative results in Tab. 1 using the inverse coefficient of variation (ICV) metric, which is a common no-reference metric used in de-striping works [36, 33, 26, 37]. We present qualitative results of de-streaked images in Fig. 6.
4.5 3D Reconstruction

### v2

@@ III-B Range/Azimuth Splatting Model
$\mu_{k}^{\prime}=K\mu_{s},\ \ \Sigma_{k}^{\prime}=J_{K}W\Sigma_{s}WJ_{K}^{T}$ (7)
where $\mu_{k}^{\prime}$ and $\Sigma_{k}^{\prime}$ are the mean and covariance transformed to range/azimuth space, $J_{K}$ is the Jacobian of $K$, and $W$ is the sensor’s view matrix, similar to the camera splatting formulation [14].
To find the intensity of acoustic returns at a given range/azimuth bin ($r_{i},\theta_{j}$), we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range/azimuth image space that overlap with the range/azimuth bin ($r_{i},\theta_{j}$). The rasterization equation to obtain the sonar image for SonarSplat is similar to that of [14], with a few key differences. SonarSplat’s rasterization equation is given by:
$I_{u}(r_{i},\theta_{j})=\frac{1}{R_{k}}\sum_{G_{k}^{\prime}\in\mathcal{G}_{i,j}}\nu_{k}~{}o_{k}~{}T_{k}~{}G_{k}^{\prime}(r_{i},\theta_{j})$ (8)
where $R_{k}$ is the range from the sensor, $T_{k}$ is the transmittance from the sensor origin to the Gaussian along the range, and $G^{\prime}_{k}$ is the Gaussian evaluated in range/azimuth space. The subscript indicates that $I_{u}$ is the unsaturated image, meaning that it does not account for the azimuth streaking that causes saturation across the image. We observe that the acoustic reflectance of a set of points depends on material properties and viewing angle, so $\nu_{k}$ uses spherical harmonics to encode view-dependent effects. In practice, our sonar images are gain adjusted, so $\frac{1}{R_{k}}$ is dropped.
III-C Azimuth Streak Modeling (ASM)
Azimuth streaking, as shown in Fig. 4, is a phenomena that is frequently observed in sonar images. It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to parallel at a specific range bin $r_{i}$ [32]. In our model, we assign each Gaussian a probability that it will cause an azimuth streak. The benefit of defining this probability per-Gaussian is that it can be optimized to produce azimuth streaks that are multi-view consistent. Let $p_{k}$ be the probability that Gaussian $k$ will contribute to an azimuth streaking phenomena. We can then splat these per-Gaussian probabilities into a range/azimuth bin to obtain an azimuth streak probability image $P_{a}$:
$P_{a}(r_{i},\theta_{j})=\sum_{G^{\prime}_{k}\in\mathcal{G}_{i,j}}p_{k}~{}o_{k}~{}T_{k}~{}G^{\prime}_{k}(r_{i},\theta_{j})$ (9)
@@ III-C Azimuth Streak Modeling (ASM)
III-C1 Adaptive Gain Module
We introduce a novel adaptive gain term, $A$, to transform the azimuth streak probability image into receiver gains. We begin designing this gain term with a few observations. First, when no Gaussian in a range bin $r_{i}$ has a high probability of azimuth streaking, the gain should be unity. Second, suppose a single Gaussian $G_{k}$ exhibits a high probability of azimuth streaking across the range bin $r_{i}$. In that case, we wish to adaptively suppress the other Gaussians in $r_{i}$ and assign a high gain to $G_{k}$. Finally, if multiple Gaussians in $r_{i}$ have high probabilities of contributing to an azimuth streak, higher gain should be assigned to those Gaussians. Following these observations, the adaptive gain term is generated for each pixel: $(r_{i},\theta_{j})$:
$A(r_{i},\theta_{j})=P_{a}(r_{i},\theta_{j})M_{a}(r_{i})\frac{e^{\gamma\cdot P_{a}(r_{i},\theta_{j})}-1}{e^{\gamma}-1}+(1-M_{a}(r_{i}))$
where $\gamma$ is a scaling factor that dictates the steepness of the adaptive gain. This behavior aligns with the insight presented in [32]: if a certain percentage of returns exceeds a threshold, an operation is applied to all of the values in a range interval. However, our proposed model differs because we do not restrict the gain to a fixed function (quadratic) but rather offer a family of curves for our splatting model to explore during optimization. The final sonar image $\hat{I}(r_{i},\theta_{j})$ with azimuth streaks is then computed by:
@@ III-C1 Adaptive Gain Module
Figure 3: Illustration of the proposed Elevation Sampling Densification Strategy (ESDS). We randomly sample $N_{p}$ pixels (shown as green sectors) based on the magnitude of $\mathcal{L}$ then randomly place $N_{g}$ Gaussians (shown as green ellipses) on the elevation arc of that pixel.
@@
Figure 4: We present our robotic platform and selected images from our diverse set of 8 sequences used in our experiments. We show RGB images in top right corners for visualization. Our evaluation datasets focus on smaller objects and larger-scale structures such as the test-tank basin. Azimuth streaks clearly occur in the infra_360 and piling_1 sequences. Images are range/azimuth.
III-D Model Optimization
We optimize for desired parameters using gradient descent and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$ [14]. Next, we note that large Gaussians cause increased training times and more distortion to the model presented in Sec. III-B. To prevent Gaussians from becoming too large, we use $\mathcal{L}_{size}=\text{ReLU}(s-s_{init})$, which encourages the scale of the Gaussians $s$ to remain smaller than their initialized scale $s_{init}$ [16]. Additionally, we find it useful for reconstruction tasks to supervise the opacities in regions with sonar returns via an opacity loss $\mathcal{L}_{o}$, similar to [35]. We compute $\mathcal{L}_{o}$ between the predicted opacity map $\hat{I}_{o}$ and $I_{\text{GT}}$ thresholded at $\tau_{o}$ to isolate strong returns. The threshold parameter is found experimentally for each dataset. $\mathcal{L}_{o}$ then becomes an $l1$ loss between $\hat{I}_{o}$ and $I^{o}_{\text{GT}}$. The final loss becomes:
$\mathcal{L}=\lambda_{l1}\mathcal{L}_{l1}+(1-\lambda_{l1})\mathcal{L}_{ssim}+\lambda_{o}\mathcal{L}_{o}+\lambda_{size}\mathcal{L}_{size}$ (13)
@@ III-D Model Optimization
where $\lambda_{l1},\lambda_{o}~{}\text{and}~{}\lambda_{size}$ are the weights for the $l1$, opacity, and size losses respectively.
III-E Elevation Sampling Densification Strategy (ESDS)
Splitting and cloning Gaussians in image space is unintuitive for imaging sonar given the elevation angle ambiguity. It is possible that a Gaussian can be split then immediately occluded by high opacity Gaussians in front of it. To address this, we introduce an alternative strategy for densification of Gaussians called the Elevation Sampling Densification Strategy (ESDS) shown in Fig. 3. Given predicted image $\hat{I}$ and ground truth $I_{\text{GT}}$, we sample $N_{p}$ pixels, where each pixel’s probability is proportional to the magnitude of the loss at that pixel. We initialize $N_{g}$ Gaussians on each elevation arc specified by the sampled pixel locations. This primarily serves to provide SonarSplat enough opportunities to resolve the elevation angle ambiguity and any possible occlusions by placing Gaussians at different elevation angles. As the optimization continues, low-opacity Gaussians are pruned similar to [14].
III-F Implementation Details
SonarSplat is built on the gsplat repo [52]. We train our method on an NVIDIA RTX 3090 GPU with 24 GB of vRAM. We calculate the frames per second (FPS) of all methods on the same NVIDIA A6000 GPU with 50 GB of vRAM. To properly optimize for the azimuth streaking probabilities $p_{k}$ of each Gaussian, we first identify where azimuth streaks occur by calculating the average intensity of the range interval $r_{i}$. Then, for the first $N_{s}$ iterations, we train only on pixels in range bins that exceed a certain average intensity threshold $\tau_{s}$. This avoids incorrectly supervising the splat in azimuth streaked regions. Importantly, $p_{k}$ is not optimized during this interval. After $N_{s}$ iterations, we optimize only $p_{k}$ in order to capture the azimuth streaking effects. We do not optimize $\mu_{k},~{}\Sigma_{k},~{}o_{k}$, or $\nu_{k}$ during this interval. This way, we allow the Gaussians to initialize and fit the images before isolating and optimizing the azimuth streaking parameters.
IV Results & Experiments
@@
Figure 5: We present qualitative novel view synthesis results from selected datasets and selected baselines. Note that SonarSplat produces more realistic sonar image synthesis compared to baselines and is better able to capture finer details in the environment. Note that ZSplat* indicates ZSplat trained only with the sonar loss. Images are Cartesian.
IV-A Robotic Platform and Datasets
We note that real, public datasets of posed sonar images suitable for novel view synthesis and reconstruction experiments do not exist. Rather, we collect sequences using a BlueRobotics BlueROV2 platform, shown in Fig. 4. Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar operating at 1.2MHz, capturing sequences between 3-20m in range, with a 130 horizontal field of view and a 20 vertical field of view. We tilt the sonar both at 30 and 0 downwards for various sequences. We use this platform to collect 8 trajectories inspecting a range of objects and in both a test tank and river environment. We use 7 sequences for evaluating image synthesis, and 1 sequence with higher coverage of the inspected object for evaluating 3D reconstruction. Sensor poses are estimated by filtering a Doppler Velocity Logger and IMU, similar to [1, 33]. We construct ground truth for 3D reconstruction using RGB images and MASt3R, scaling the reconstruction to metric scale [17]. We find that MASt3R produces less artifacts than traditional photogrammetry pipelines like AgiSoft Metashape [21].
IV-B Baselines
For baselines, we compare to Neusis [33], Neusis-NGP [50], DSC [6], and ZSplat* [35]. Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with only the sonar loss since there is no other work that has explored Gaussian splatting for imaging sonar [35]. All baselines were trained using official released code and hyperparameters were tuned to the best of our ability.
TABLE I: Novel view synthesis performance on real world held-out validation data. Note that ZSplat* is ZSplat with only the sonar loss activated for training. First, second, and third best are shown in red, orange, and yellow, respectively.
@@ IV-B Baselines
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
Neusis [33] 14.68 0.04 0.64 0.03
Neusis-NGP [50] 19.56 0.33 0.64 0.01
DSC [6] 14.70 0.07 0.66 0.03
ZSplat* [35] 10.06 0.02 0.69 155.32
SonarSplat (ours) 22.79 0.41 0.48 13.45
IV-C Novel View Synthesis
We report validation PSNR, SSIM, and LPIPS following [14], using the lpipsPyTorch library. For our validation set, we take every 8th image from the dataset as introduced in [26]. We report SonarSplat’s performance compared to baselines in Tab. I. We find that SonarSplat consistently synthesizes more realistic sonar images, quantified by its higher PSNR and SSIM, and lower LPIPS values. Qualitative results from selected sequences are presented in Fig. 5. We also report the rendering speed in FPS of each method in Tab. I. SonarSplat and ZSplat* both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods. ZSplat* is optimized for high rendering speed, which will be a focus of future work.
SonarSplat Rendering Ablations
VDR
ASM
ESDS
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
21.32 0.35 0.53
21.16 0.34 0.55
22.51 0.40 0.52
22.79 0.41 0.48
TABLE II: Ablation study results for the proposed acoustic reflectance parameter (VDR), azimuth streak modeling (ASM), and elevation sampling densification strategy (ESDS). Best in bold.
IV-D Ablation Studies
We perform ablation studies on the view-dependent reflectivity parameter $\nu_{k}$ (VDR), the azimuth streak modeling (ASM), and the Elevation Sampling Densification Strategy (ESDS). Ablation studies are averaged across 7 evaluation datasets and trained with exactly the same hyperparameters as results from Sec. IV-C. Methods that don’t use ESDS use the mcmc strategy in gsplat [52]. Results are reported in Tab. II and shown in Fig. 6. Clearly, the addition of azimuth streak modeling, view-dependent acoustic reflectivity, and the elevation sampling densification strategy improves the image synthesis capabilities of SonarSplat.
Figure 6: Qualitative comparison of SonarSplat without and with azimuth streak modeling (ASM) for piling_1. Note that if ASM is not modeled, rendered views will produce erroneous streaks (indicated with red arrows) instead of capturing the true streaks (indicated with blue arrows). Images are polar.
Figure 7: Qualitative results for azimuth streak removal capabilities of SonarSplat. Locations of azimuth streaks in the captured images are indicated by blue arrows.
TABLE III: Quantitative results for azimuth streak removal. We report the inverse coefficient of variation (ICV). Higher is better ($\uparrow$) and best is in bold.
De-Streaking Performance (ICV) $\uparrow$
Captured Image De-Streaked Prediction
Average 2.91 18.58
IV-E Azimuth Streak Removal/De-Streaking
@@
Using SonarSplat, we can undo the adaptive gain to render images without azimuth streaks present and recover suppressed returns. This vision task is most similar to image de-striping from the remote sensing literature. We present quantitative results in Tab. III using the inverse coefficient of variation (ICV) metric, which is a common no-reference metric used in de-striping works [42, 36, 29]. We present qualitative results of de-streaked images in Fig. 7. The Azimuth Streak Modeling allows SonarSplat to qualitatively and quantitatively remove streaks effectively and improves the finer features in sonar images.
Figure 8: Qualitative results from 3D reconstruction of SonarSplat compared to baselines. We inspect a concrete piling in a test tank. Ground truth is created using collected RGB images.
IV-F Real-World 3D Reconstruction
3D reconstruction is crucial for inspection and visualization of underwater structures. In a test tank, we demonstrate inspection of a concrete piling, a structure commonly found in bridges and dams.